The decomposition of NO2(g) : 2NO2(g) ----->2NO(g) +O(g) is a second order reaction and the rate constant is 0.755L/(mol-s). In an experiment, the initial concentration of NO2(g) was 0.00650 mol/L. (a)What is the concentration of NO(g) after 125 seconds have elapsed? (b) How many seconds will it take for the concentration of NO2(g) to drop to 0.00100 mol/L?


1
Expert's answer
2021-03-25T04:10:02-0400

(a) For second order reaction:

K = 1/t ( 1/At - 1/A0)

0.755 = 1/125 ( 1/At - 1/0.065)

94.375 = 1/At - 1/0.065

248.221 = 1/At

At = 0.00402 M

(b) 0.755 = 1/t ( 1/0.00100- 1/0.0065)

0.755 = 1/t (1000 - 153.84)

0.755 = 846.16/t

t = 1120s